OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (6, 2) and has a slope of 3? y = ------ x + ------

OpenStudy (akashdeepdeb):

The slope-intercept form of a line is: \(y = mx + c\) where c is the y-intercept of the line. You are given a point on the line (6,2) and you are given the slope of the line (3). The point-slope form of a line is: \(y - y_1 = m(x - x_1)\) because \(m = \frac{y-y_1}{x-x_1}\) where m is the slope of the line, \((x_1,y_1)\) is the point on the line. After substituting the values represent it in the slope-intercept form.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

uhm i'm really bad at math so is there any way you could help me simplify that answer because i'm lost....

OpenStudy (anonymous):

y−y1=m(x−x1) <-- use this equation. You have 2 points (x,y) = (6,2) and m which is the slope m = 3.

OpenStudy (anonymous):

just plug in and solve for y

OpenStudy (anonymous):

thank you! :)

OpenStudy (anonymous):

was you able to do it?